OpenStudy (anonymous):

A chemist need 190 milliliters of a 41% solution but has only 33% and 53% solutions available. Find how many milliliters of each should be mixed to get the desired solution.

OpenStudy (btaylor):

Let's name the 33% solution A, and the 51% solution B. We will make a system: .33A + .53B = .41(190) --> 33A + 53B = 7790 A + B = 190 The first one gives the concentration of each element. The second gives the total volume. Solving using subraction: 33A + 53B = 7790 -33A -33B = -6270 0 + 20B = 1520 B = 76 --> A = 190-76=114 So you need 114 mL A and 76 ml B.

OpenStudy (anonymous):

okay thanks so do i just supstitute my numbers in that equation to find the answer?